The probability that an individual without college education earns more than $30000 is 0.2, where as the probability that a person with B.S. or higher degree earns $30000 is 0.6. the probability that a person chosen at random has B.S. degree is 0.3. what is the probability that a person has atleast a B.S. degree if it is known that he/she earns more than $30000?.
1
Expert's answer
2020-06-29T18:49:07-0400

Formula:


Baye's formula for trees:


P(E/F)= P(E and F)/ P(F)


Hence to find the probability that a person has atleast a B.S. degree if it is known that he/she earns more than $30000=

P(at least BS degree or higher and earn more than $30000)/P(more than $30000)




From the tree diagram ,

= (0.3)*(0.6)/(0.6+0.2)

=0.18/0.8

=0.225